# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 4.3 Randomized ID/CUR
In both the randomized rangefinder and the randomized SVD, orthonormal matrices are computed whose columns form an approximate basis for the row or column space of an input matrix $\bm{\mathsf{A}}$. The entries of these orthonormal matrices are linear combinations of the entries of $\bm{\mathsf{A}}$. However, for some problems (e.g., if $\bm{\mathsf{A}}$ is sparse or non-negative), it may be preferable to have a “natural basis,” consisting of subsets of rows or columns of $\bm{\mathsf{A}}$, as in the ID/CUR factorizations of Sect. 2.4. In this section, we provide an overview of randomized methods to compute approximate ID/CUR factorizations based on the randomized rangefinder, including recent work to improve skeleton selection and performance DongRBRP2025; dong2022; Pearce2025; Pritchard2025.
With the randomized rangefinder
Recall from Sect. 2.4 that pivoting algorithms, such as column-pivoted QR (CPQR) or LU with partial pivoting, are frequently used in the deterministic setting to select skeletons for an ID/CUR factorization. CPQR produces skeletons that are close to optimal in practice Cheng2005; martinsson2006interpolation. In particular, this is true when CPQR is applied to a random sketch of $\bm{\mathsf{A}}$ with appropriate embedding dimension. It is also conjectured in (trefethen1990, Remark 4) and exhibited empirically in dong2022; Pearce2025 that randomized sketching equips LUPP with near-optimal skeleton selection capability as well, even in adversarial cases. We illustrate how randomized sketching with CPQR or LUPP can be used to compute a row ID, then discuss alternate skeleton selection strategies that are rank-adaptive or based on sampling instead of sketching.
More specifically, let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be an input matrix of exact rank $k$ for notational convenience. Let $\bm{\mathsf{Y}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ have columns that span the range of $\bm{\mathsf{A}}$, e.g., the output of the rangefinder algorithm. By definition,
$\bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{Y}}}\ \underset{k\times n}{\bm{\mathsf{F}}}$
for some matrix $\bm{\mathsf{F}}$. Compute a row ID of $\bm{\mathsf{Y}}$, either using CPQR via $\sim,\sim,I]=\texttt{qr}(\bm{\mathsf{Y}}^{*},k)$ as in Sect. 2.4 or LUPP via $\sim,\sim,I]=\texttt{lu}(\bm{\mathsf{Y}},k)$ as in Sect. 2.4, to obtain
$\bm{\mathsf{Y}}=\underset{m\times k}{\bm{\mathsf{X}}}\underset{k\times k}{\bm{\mathsf{Y}}(I,:)}.$ (42)
We now observe that index set $I$ and row interpolation matrix $\bm{\mathsf{X}}$ form a row ID of $\bm{\mathsf{A}}$:
@@
Thus, to compute a row ID of $\bm{\mathsf{A}}$, it suffices to compute a row ID of a matrix $\bm{\mathsf{Y}}$ whose columns span the range of $\bm{\mathsf{A}}$.
We can construct such a matrix $\bm{\mathsf{Y}}\in\mathbb{F}^{m\times\ell}$ for a rank-$k$ approximation of $\bm{\mathsf{A}}$ using the randomized rangefinder with high probability with, e.g., $\ell=k+10.$ Additionally, since we do not need $\bm{\mathsf{Y}}$ to be an orthonormal matrix here, we can save on computational cost by observing that the range of $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is itself an accurate approximation to the range of $\bm{\mathsf{A}}$ with high probability (halko2011, Theorems 10.5 and 10.6). The randomized rangefinder error $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ using $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is equal to $\|\bm{\mathsf{A}}-\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}\bm{\mathsf{A}}\|$, where $\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}$ is the orthogonal projector onto the column space of $\bm{\mathsf{Y}}$ or, equivalently, the column space of $\bm{\mathsf{Q}}$. Pseudocode for this randomized row ID algorithm (with sketched CPQR) is provided in Algorithm 4.
@@
Adaptive algorithms for randomized ID/CUR
Frequently in applications, the target rank for a low-rank approximation is unknown a priori, in which case we turn to rank-adaptive methods. If using CPQR for skeleton selection, the adaptive version of the randomized rangefinder may be used to build an orthonormal basis incrementally from independent random sketches or samples.
However, if we are using LUPP as the skeleton selection method for improved efficiency, we want to estimate the approximation error without computing an orthonormal basis. We briefly illustrate how a row ID can be constructed with the rank-adaptive, LUPP-based method of Pearce2025. Given an error tolerance $\tau$ and input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{m\times n}$, after $t$ iterations of this algorithm, we obtain target rank $k_{t}$, row skeletons $I_{s}^{(t)}$ with $|I_{s}^{(t)}|=k_{t}$, and interpolation matrix $\bm{\mathsf{W}}^{(t)}$ such that
@@
A similar idea is presented in Pritchard2025 for a rank-adaptive CUR factorization, using a single small sketch of the residual matrix $\bm{\mathsf{S}}=\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{U}}\bm{\mathsf{R}}$. (Here, $\bm{\mathsf{U}}$ is computed as $\bm{\mathsf{A}}(I,J)^{\dagger}$ vs. $\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$ as in Sect. 2.4.) The column residual is initialized as a row sketch, $\bm{\mathsf{S}}^{(C)}_{0}=\bm{\mathsf{G}}\bm{\mathsf{A}}\in\mathbb{F}^{1.1b\times n}$, where typically $5\leq b\leq 250$, and the row and column skeletons are empty sets. While the error criterion is unmet, column skeletons $J$ are first computed from $\bm{\mathsf{S}}^{(C)}_{0}$, and the row residual is formed as $\bm{\mathsf{S}}^{(R)}_{1}=\bm{\mathsf{A}}(:,J)-\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}(:,J)=\bm{\mathsf{A}}(:,J)$. Row skeletons $I$ are then computed from $\bm{\mathsf{S}}^{(R)}_{1}$, and the column residual is formed as $\bm{\mathsf{S}}^{(C)}_{1}=\bm{\mathsf{G}}\bm{\mathsf{A}}-\bm{\mathsf{G}}\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}$, where now $\bm{\mathsf{C}}_{1}=\bm{\mathsf{A}}(:,J)$, $\bm{\mathsf{R}}_{1}=\bm{\mathsf{A}}(I,:)$, and $\bm{\mathsf{U}}_{1}=\bm{\mathsf{A}}(I,J)^{\dagger}$ are no longer empty. This procedure is repeated, with new skeletons appended to previously selected ones, until the desired approximation accuracy is achieved.
Algorithms based on coordinate sampling
@@
In ChenEpperlyTroppWebber2025, the randomly pivoted (RP) Cholesky algorithm is introduced for positive semi-definite matrices $\bm{\mathsf{A}}\in\mathbb{F}^{m\times m}$. Given a target rank $k$, the algorithm outputs a set of pivot indices $S=\{s_{1},\ldots,s_{k}\}$ akin to skeletons, and a matrix $\bm{\mathsf{F}}\in\mathbb{F}^{m\times k}$ that defines the Nyström approximation $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{F}}\bm{\mathsf{F}}^{*}$. We begin by defining $\bm{\mathsf{d}}=\textup{diag}(\bm{\mathsf{A}})$ and initializing the matrix $\bm{\mathsf{F}}$ as $\bm{\mathsf{0}}$.
After sampling each pivot $s_{i}\sim\bm{\mathsf{d}}/\sum_{j=1}^{m}\bm{\mathsf{d}}(j)$, we form $\bm{\mathsf{g}}=\bm{\mathsf{A}}(:,s_{i})$ and remove any overlap with previously selected columns via
@@ 5 Tensor Preliminaries
This section marks the beginning of our exploration of tensors. We begin with formal notation and definitions in Sect. 5.1, presenting tensors as natural higher-order generalizations of matrices, from 2-way to $d$-way arrays. As a result, many of the randomized methods in the previous sections developed for matrices can be extended to tensors, though there are subtleties with higher-order inputs, and many concepts can be obscured by the unwieldy notation that frequently accompanies them. We try to mitigate this by providing visualizations, including a table of notation (Table 4) and illustrations of the matrix and tensor products in Sect. 5.2. In Sect. 5.3, we point out some peculiarities specific to higher-order tensors, such as tensor rank and uniqueness of decompositions.
While there are many different ways to represent tensors, we focus on two fundamental tensor decomposition formats: the canonical polyadic (CP) and Tucker. In Sect. 5.3, we lay out key concepts and practicalities for each of them, including algorithms commonly used to compute them. We then discuss particular low-rank tensor decompositions that can be represented in each format, such as higher-order extensions of the matrix SVD and ID/CUR. Throughout this section, we try to emphasize that the classical matrix methods covered in previous sections underlie those for tensors.
Symbol Definition
$\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$ Tensor with $d$ modes
$\|\bm{\mathcal{X}}\|$ Frobenius norm $\left(\sum_{i_{1}=1}^{N_{1}}\ldots\sum_{i_{d}=1}^{N_{d}}\bm{\mathcal{X}}(i_{1},\ldots,i_{d})^{2}\right)^{1/2}$
$N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$ Number of entries of $\bm{\mathcal{X}}$
$N^{(-j)}=\prod_{k\neq j}N_{j}$ Number of entries of $\bm{\mathcal{X}}$ excluding mode $j$
$\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$ Mode-$j$ unfolding of $\bm{\mathcal{X}}$
$\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}$ Kronecker product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{K\times L}$
$\bm{\mathsf{A}}\odot\bm{\mathsf{B}}$ Khatri-Rao product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times K},\bm{\mathsf{B}}\in\mathbb{F}^{J\times K}$
$\bm{\mathsf{A}}*\bm{\mathsf{B}}$ Hadamard product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$
Table 4: Summary of tensor notation.
@@ 5.1 Notation and Definitions
We define $N^{(-j)}=\prod_{k=1,k\neq j}^{d}N_{k}$ as the product of all dimensions except mode $j$, and $N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$. Through a rearrangement of entries known as matricization, a $d$-mode tensor can be “unfolded” into a matrix. The mode-$j$ unfolding of a $d$-mode tensor $\bm{\mathcal{X}}$, denoted $\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$, $j=1,\ldots,d$, is a matrix whose columns are the mode-$j$ fibers of $\bm{\mathcal{X}}$. Throughout this section, we assume that the $d$ modes are ordered $1,\ldots,d$ for notational simplicity, but for computational reasons, it may be beneficial to re-order the modes for processing.
@@ 5.2 Matrix and Tensor Products
The Kronecker product then takes the form
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}\otimes\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,L}\\ \vdots&\ddots&\vdots\\ b_{K,1}&\cdots&b_{K,L}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,1}b_{1,L}&\cdots\cdots&a_{1,J}b_{1,1}&\cdots&a_{1,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{1,1}b_{K,1}&\cdots&a_{1,1}b_{K,L}&\cdots\cdots&a_{1,J}b_{K,1}&\cdots&a_{1,J}b_{K,L}\\ \vdots&&&&&&\vdots\\ \vdots&&&\ddots&&&\vdots\\ \vdots&&&&&&\vdots\\ a_{I,1}b_{1,1}&\cdots&a_{I,1}b_{1,L}&\cdots\cdots&a_{I,J}b_{1,1}&\cdots&a_{I,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{I,1}b_{K,1}&\cdots&a_{I,1}b_{K,L}&\cdots\cdots&a_{I,J}b_{K,1}&\cdots&a_{I,J}b_{K,L}\end{bmatrix}}$
$=\begin{bmatrix}a_{1,1}\bm{\mathsf{B}}&\cdots&a_{1,J}\bm{\mathsf{B}}\\ \vdots&\ddots&\vdots\\ a_{I,1}\bm{\mathsf{B}}&\cdots&a_{I,J}\bm{\mathsf{B}}\end{bmatrix}$.
@@
for $k=1,\ldots,K.$ The Khatri-Rao product takes the form
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,K}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,K}\end{bmatrix}}\odot\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,K}\\ \vdots&\ddots&\vdots\\ b_{J,1}&\cdots&b_{J,K}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\odot\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&a_{1,2}b_{1,2}&\cdots&a_{1,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,1}b_{J,1}&a_{1,2}b_{J,2}&\cdots&a_{1,K}b_{J,K}\\ \vdots&\vdots&&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \vdots&\vdots&&\vdots\\ a_{I,1}b_{1,1}&a_{I,2}b_{1,2}&\cdots&a_{I,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{I,1}b_{J,1}&a_{I,2}b_{J,2}&\cdots&a_{I,K}b_{J,K}\\ \end{bmatrix}}$
$=\begin{bmatrix}\bm{\mathsf{a}}_{1}\otimes\bm{\mathsf{b}}_{1}&\bm{\mathsf{a}}_{2}\otimes\bm{\mathsf{b}}_{2}&\cdots&\bm{\mathsf{a}}_{K}\otimes\bm{\mathsf{b}}_{K}\end{bmatrix}$,
where $\bm{\mathsf{a}}_{i}$ and $\bm{\mathsf{b}}_{i}$ are the columns of $\bm{\mathsf{A}}$ and $\bm{\mathsf{B}}$. There is a useful property relating the Kronecker and Khatri-Rao products: if $\bm{\mathsf{A}}\in\mathbb{F}^{K\times J}$, $\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$, $\bm{\mathsf{C}}\in\mathbb{F}^{I\times K}$, and $\bm{\mathsf{D}}\in\mathbb{F}^{J\times I}$,
$(\bm{\mathsf{C}}\otimes\bm{\mathsf{D}})(\bm{\mathsf{A}}\odot\bm{\mathsf{B}})=(\bm{\mathsf{C}}\bm{\mathsf{A}})\odot(\bm{\mathsf{D}}\bm{\mathsf{B}}).$
Additionally we have the following inequality of matrix norms:
$\|\bm{\mathsf{A}}\odot\bm{\mathsf{B}}\|^{2}_{\text{F}}\leq\|\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}\|^{2}=\|\bm{\mathsf{A}}\|^{2}\|\bm{\mathsf{B}}\|^{2}.$
@@
The Hadamard product can be visualized as element-wise multiplication,
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}*\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,J}\\ \vdots&\ddots&\vdots\\ b_{I,1}&\cdots&b_{I,J}\end{bmatrix}}=\underset{\bm{\mathsf{A}}*\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,J}b_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}b_{I,1}&\cdots&a_{I,J}b_{I,J}\end{bmatrix}}.$
@@
$\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}_{j}\bm{\mathsf{X}}_{(j)}\left(\bm{\mathsf{U}}_{d}\otimes\cdots\otimes\bm{\mathsf{U}}_{j+1}\otimes\bm{\mathsf{U}}_{j-1}\otimes\cdots\otimes\bm{\mathsf{U}}_{1}\right)^{\top}.$

### v2

@@ 4.3 Randomized ID/CUR
In both the randomized rangefinder and the randomized SVD, orthonormal matrices are computed whose columns form an approximate basis for the row or column space of an input matrix $\bm{\mathsf{A}}$. However, for some problems (e.g., if $\bm{\mathsf{A}}$ is sparse or non-negative), it may be preferable to have a “natural basis,” consisting of rows or columns of $\bm{\mathsf{A}}$, as in the ID/CUR of Sect. 2.5. In this section, we provide an overview of randomized methods to compute approximate ID/CUR factorizations, including recent work to improve skeleton selection and performance DongRBRP2025; dong2022; Pearce2025; Pritchard2025. We first illustrate how a random sketch can be used as a proxy for the input matrix in ID/CUR approximations.
Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be of exact rank $k$ for notational convenience. Let $\bm{\mathsf{Y}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ have columns that span the range of $\bm{\mathsf{A}}$, not necessarily orthogonal. By definition, $\bm{\mathsf{A}}={\bm{\mathsf{Y}}}{\bm{\mathsf{F}}}$ for some matrix $\bm{\mathsf{F}}\in\mathbb{F}^{k\times n}$. Compute a row ID of $\bm{\mathsf{Y}}$, e.g., with CPQR via $\sim,\sim,I]=\texttt{qr}(\bm{\mathsf{Y}}^{*},k)$ or LUPP via $\sim,\sim,I]=\texttt{lu}(\bm{\mathsf{Y}},k)$, so that $\bm{\mathsf{Y}}=\bm{\mathsf{X}}\bm{\mathsf{Y}}(I,:)$. We now observe that index set $I$ and row interpolation matrix $\bm{\mathsf{X}}$ form a row ID of $\bm{\mathsf{A}}$:
@@
Thus, to compute a row ID of $\bm{\mathsf{A}}$, it suffices to compute a row ID of a matrix $\bm{\mathsf{Y}}$ whose columns span the range of $\bm{\mathsf{A}}$, and since we do not need $\bm{\mathsf{Y}}$ to be orthogonal, we can save on computational cost by using the fact that the range of $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is itself an accurate approximation to the range of $\bm{\mathsf{A}}$ with high probability (halko2011, Theorems 10.5 and 10.6). We observe that the randomized rangefinder error $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ using $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is equal to $\|\bm{\mathsf{A}}-\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}\bm{\mathsf{A}}\|$, where $\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}$ is the orthogonal projector onto the column space of $\bm{\mathsf{Y}}$ (or, equivalently, the column space of $\bm{\mathsf{Q}}$). Pseudocode for this randomized row ID procedure (with sketched CPQR) is provided in Algorithm 4.
@@
With rank-adaptive randomized sketching and greedy pivoting
Frequently in applications, the target rank for a low-rank approximation is unknown a priori, in which case we turn to rank-adaptive methods. If using CPQR for skeleton selection, the adaptive version of the randomized rangefinder may be used to build an orthonormal basis incrementally from independent random sketches or samples. However, if we are using LUPP as the skeleton selection method for improved efficiency and better parallelizability, we want to estimate the approximation error without computing an orthonormal basis. We briefly illustrate how a row ID can be constructed with the rank-adaptive, LUPP-based method of Pearce2025. Given an error tolerance $\tau$ and input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{m\times n}$, after $t$ iterations of this algorithm, we obtain target rank $k_{t}$, row skeletons $I_{s}^{(t)}$ with $|I_{s}^{(t)}|=k_{t}$, and interpolation matrix $\bm{\mathsf{W}}^{(t)}$ such that
@@
A similar idea is presented in Pritchard2025 for a rank-adaptive CUR factorization, using a single small sketch of the residual matrix $\bm{\mathsf{S}}=\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{U}}\bm{\mathsf{R}}$. (Here, $\bm{\mathsf{U}}$ is computed as $\bm{\mathsf{A}}(I,J)^{\dagger}$ vs. $\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$ as in Sect. 2.5.) The column residual is initialized as a row sketch, $\bm{\mathsf{S}}^{(C)}_{0}=\bm{\mathsf{G}}\bm{\mathsf{A}}\in\mathbb{F}^{1.1b\times n}$, where typically $5\leq b\leq 250$, and the row and column skeletons are empty sets. While the error criterion is unmet, column skeletons $J$ are first computed from $\bm{\mathsf{S}}^{(C)}_{0}$, and the row residual is formed as $\bm{\mathsf{S}}^{(R)}_{1}=\bm{\mathsf{A}}(:,J)-\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}(:,J)=\bm{\mathsf{A}}(:,J)$. Row skeletons $I$ are then computed from $\bm{\mathsf{S}}^{(R)}_{1}$, and the column residual is formed as $\bm{\mathsf{S}}^{(C)}_{1}=\bm{\mathsf{G}}\bm{\mathsf{A}}-\bm{\mathsf{G}}\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}$, where now $\bm{\mathsf{C}}_{1}=\bm{\mathsf{A}}(:,J)$, $\bm{\mathsf{R}}_{1}=\bm{\mathsf{A}}(I,:)$, and $\bm{\mathsf{U}}_{1}=\bm{\mathsf{A}}(I,J)^{\dagger}$ are no longer empty. This procedure is repeated, with new skeletons appended to previously selected ones, until the desired approximation accuracy is achieved.
With rank-adaptive random pivoting
@@
In ChenEpperlyTroppWebber2025, the randomly pivoted (RP) Cholesky algorithm is introduced for positive semi-definite matrices $\bm{\mathsf{A}}\in\mathbb{F}^{m\times m}$. Given a target rank $k$, the algorithm outputs a set of pivot indices $S=\{s_{1},\ldots,s_{k}\}$ akin to skeletons, and a matrix $\bm{\mathsf{F}}\in\mathbb{F}^{m\times k}$ for approximate Cholesky factorization $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{F}}\bm{\mathsf{F}}^{*}$, which is the factorized form of the column Nystr”om approximation $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{A}}(:,S)\bm{\mathsf{A}}(S,S)^{{\dagger}}\bm{\mathsf{A}}(S,:)$. By selecting pivots, or skeletons, with probability proportional to the magnitudes of the diagonal entries of the residual matrix in each iteration, the RP Cholesky algorithm returns the Cholesky-factorized form of the Nyström approximation in $O(km)$ matrix entry evaluations and $O(k^{2}m)$ additional arithmetic operations, with $O(km)$ storage requirements.
We begin by defining $\bm{\mathsf{d}}=\textup{diag}(\bm{\mathsf{A}})$ and initializing the matrix $\bm{\mathsf{F}}$ as $\bm{\mathsf{0}}$. After sampling each pivot $s_{i}\sim\bm{\mathsf{d}}/\sum_{j=1}^{m}\bm{\mathsf{d}}(j)$, we form $\bm{\mathsf{g}}=\bm{\mathsf{A}}(:,s_{i})$ and remove any overlap with previously selected columns via
@@ 5 Tensor Preliminaries
This section marks the beginning of our exploration of tensors, where we present tensors as natural higher-order generalizations of matrices, from 2-way to $d$-way arrays. We begin with formal notation and definitions in Sect. 5.1 and discuss important matrix and tensor operations in Sect. 5.2. While there are many different ways to represent tensors, we focus on two fundamental tensor decomposition formats: the canonical polyadic (CP) and Tucker. In Sect. 5.3, we lay out key concepts and practicalities for each of them, including algorithms to compute them. We then discuss particular low-rank tensor decompositions that can be represented in each format, such as higher-order extensions of the matrix SVD and ID/CUR. Throughout this section, we show how the classical matrix methods of Sect. 2 form the foundation of those for tensors.
@@ 5.1 Notation and Definitions
We define $N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$, and $N^{(-j)}=\prod_{k=1,k\neq j}^{d}N_{k}=N/N_{j}$. Through a rearrangement of entries known as matricization, a $d$-mode tensor can be “unfolded” into a matrix. The mode-$j$ unfolding of a $d$-mode tensor $\bm{\mathcal{X}}$, denoted $\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$, $j=1,\ldots,d$, is a matrix whose columns are the mode-$j$ fibers of $\bm{\mathcal{X}}$. We assume the $d$ modes are ordered $1,\ldots,d$ for simplicity, but for computational reasons, it may be beneficial to re-order the modes for processing.
@@ 5.2 Matrix and Tensor Products
for $k=1,\ldots,K.$
@@
$\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}_{j}\bm{\mathsf{X}}_{(j)}\left(\bm{\mathsf{U}}_{d}\otimes\cdots\otimes\bm{\mathsf{U}}_{j+1}\otimes\bm{\mathsf{U}}_{j-1}\otimes\cdots\otimes\bm{\mathsf{U}}_{1}\right)^{*}.$